OpenStudy (anonymous):

Choose the correct simplification of (3a2 + 8a − 4) + (5a2 − 4).

OpenStudy (anonymous):

@Chelsea04

OpenStudy (anonymous):

3a^2+5a^2=8a^2 8a=8a -4-4=-8 put these all together and get 8a^2+8a-8

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

Choose the correct simplification of (3x3 + 2x2 − 5x) − (8x3 − 2x2).

OpenStudy (anonymous):

@Chelsea04

OpenStudy (anonymous):

3x2-8x3= 2x2+2x2= -5x= try it yourself

OpenStudy (anonymous):

3x3-8x3 is the first one sorry, typo

OpenStudy (anonymous):

do i times it or minus?

OpenStudy (anonymous):

you do 3-8 and then add the x3 at the end

OpenStudy (anonymous):

-5x3

OpenStudy (anonymous):

yea :) now try the second one

OpenStudy (anonymous):

can u write it out?

OpenStudy (anonymous):

@Chelsea04

OpenStudy (anonymous):

2x2+2x2= 2+2 x2

OpenStudy (anonymous):

4x?

OpenStudy (anonymous):

4x2

OpenStudy (anonymous):

so the full answer is: -5x3+4x2-5x yea?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np